If josh can paint a house in one hour, and Noah can paint the same house in three hours. How long will it take them to paint the house together? Justify your answer. (This is the exact question we were asked, I need a formula for it please)
Answer by jim_thompson5910(35256)

1/(josh's time) + 1/(noah's time) = 1/(total time)


1/1 + 1/3 = 1/x


3/3 + 1/3 = 1/x


(3+1)/3 = 1/x


4/3 = 1/x


3/4 = x


x = 3/4


Doing the job together will take them 3/4 hours (in other words, 45 minutes)